Listed below are student evaluation ratings of​ courses, where a rating of 5 is for​ "excellent." The ratings were obtained at one university in a state. Construct a confidence interval using a 95​% confidence level. What does the confidence interval tell about the population of all college students in the​ state?

3.7​, 2.9​, 3.9​, 4.7​, 2.9​, 3.9​, 3.8​, 4.9​, 4.3​, 4.1​, 4.2​, 4.0​, 3.4​, 3.9​, 3.6
Answer #1

From the given data,
mean = 3.9
s = 0.5557 , n = 15

The t value at 95% confidence interval is,

alpha = 1 - 0.95 = 0.05
alpha/2 = 0.05/2 = 0.025

df =15 -1 =14
t(alpha/2,df) = t(0.025,14) = 2.145

t value at 95% = 2.145

Margin of error = E =z *(s/sqrt(n))
= 2.145 * (0.5557/sqrt(15))
= 0.3078

CI = mean +/- E
= 3.9 +/- 0.3078
= (3.5722 , 4.1878)

THe 95% CI is (3.5722 , 4.1878)